# The PAU Survey: galaxy stellar population properties estimate with narrow-band data

### v1

@@ 4.3 Results with Prospector
Figure 5: Hexbin plots showing the ground truth from the Prospector-$\alpha$ stellar population simulations on the x-axis versus the MAP value from t Prospector, for stellar mass $M_{*}$, current SFR, mean mass-weighted stellar ages $\langle t_{*}\rangle_{m}$, stellar metallicity $Z_{*}$ and the optical depths of the two-component dust attenuation model $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$.
@@
We find robust estimates for stellar masses and SFRs, which in this case is expected due to the same star formation histories used for the generation of the data and the forward-modelling of the fitted SED. Dust parameters are also recovered accurately on a 0.5 dex level for the majority of the sample galaxies, albeit with some objects where the birth cloud dust attenuation is not reliably recovered. The overall intrinsic scatter from the 1:1 relation apparent for every free parameter is a consequence of fits where the MCMC chain is not fully converged yet after the set number of iterations, as well as parameter degeneracies and the simulated uncertainties of the model photometry.
Our setup however struggles to estimate accurate stellar metallicity posteriors, mainly for older populations with high $Z_{*}$, where the estimated values seem to be mostly clipped to the edges of the uniform prior space. We attribute this to the prior differences during the simulation and the sampling phase, where we once chose a joint mass-metallicity prior and later a uniform distribution for sampling. Increasing the number of likelihood calls with more MCMC iterations or more walkers can help to alleviate this problem, as we restricted these hyper-parameters due to computational limitations. We show in App. A an example fit with an increased iteration count, which can lead to substantial shift of the best-fit parameters. Additionally, most likely due to the age-metallicity degeneracy, $\langle t_{*}\rangle_{m}$ is not estimated well for a substantial subset of mock objects, where especially the ages of young stellar populations are overestimated by up to one order of magnitude. This then also influences the metallicity estimates and hinders an accurate measurement of both of these parameters without a joint prior.
Figure 6: Distribution of MAP values of the selected properties as determined by cigale: stellar mass, mass-weighted stellar age, SFR over 100 Myrs, stellar metallicity and optical depths of ISM and BC dust components. Blue shows the results from using narrow-band and broad-band data, red depicts the properties from using CFHTLS $ugriz$ alone.
@@
Additionally, we note that the reliance on only the MAP value of the posterior distribution is usually not always sufficient, mostly due to degeneracies of the parameters. Looking at some full posteriors from our estimates, one can identify multiple peaks with a high likelihood that where found exploring different regions of a joint dust-age-metallicity prior. We show an example of such a posterior distribution in App. A.
@@ 5 Stellar population properties
Using our selection of fitting hyper-parameters and cigale grid values, we show the results from both codes on the sample of 25 491 galaxies in the COSMOS field. We complete two runs for cigale on the catalog, once using CFHTLS $ugriz$ data only and once with additional 40 PAUS narrow-band fluxes. For Prospector, we only run the code once with narrow-band+broad-band data, due to the required $\sim$ 500k CPU hours for one run on the full sample. We first show the overall distributions of our parameter estimation as well as the robustness of the inferred properties. We also perform a comparison between the two codes and the COSMOS2020 catalog to determine the predictive power of the PAUS narrow-band filter set in relation to previous studies with broad filter baselines that also include narrow-band photometries.
@@ 5.1 SED fitting with CIGALE
In Fig. 6 we show the distributions of the inferred physical properties obtained with cigale on our catalogue of 25 491 galaxies in the COSMOS field. The blue contours and histograms show the distributions resulting from running the SED fitting routine on the combined PAUS (narrow-band) and CFHTLS (broad-band) dataset, while the red contours were determined by using the broad-band $ugriz$ data alone.
We first note that the stellar mass, which usually represents the most robust property inferred by SED fitting, is similarly constrained by either only the broad-bands or the combination of narrow-bands and broad-band data. The other properties, especially metallicity and dust component optical depths, show that with the addition of narrow-bands we can capture a broader range of physical values with respect to the sole use of broad-bands. Since these physical properties are strongly correlated to distinct spectral features, the net effect of an increased photometric resolution manifests in the form of a more diverse distribution. Broad-band filters alone tend to estimate metallicities that only slightly differ from $Z_{*}=Z_{\odot}$, as well as ages towards the high-end tail of the mass-weighted age distribution and low optical depths $\hat{\tau}_{\lambda,2;1}$ of the dust attenuation model. This can be interpreted as a result of the large impact of the prior choice and size on these parameters in the broad-band only case, resulting in a narrower posterior.
Nevertheless, the effect of the narrow-band data is not as strong as expected, where most of the sample galaxies are still similarly fitted with reasonable property estimates from the underlying SPS forward-model using $ugriz$ data alone. However, we see that narrow-bands do not lead to to a different region of the parameter space with respect to the use of broad-bands only, which indicates that the model is performing in the same manner with either setup. Additionally, we note that the overall coverage of the prior ranges set by the cigale parameter space from Tab. 2 is largely diminished in the posterior distribution. This means that either the priors are too wide and the true observations do not cover the ranges set from the Prospector-$\alpha$ model, or the sparse grid leads to an incomplete sampling of the high-dimensional parameter space. This is facilitated by the degeneracies of certain model properties. We fill further explore this after analyzing the Prospector results and the comparison with COSMOS2020.
@@
with $S_{0}$, $M_{0}$, $\alpha$ and $\beta$ as empirical parameters that depend on redshift. Popesso et al. (2023) recently presented a new parametrization of the star forming main sequence as a function of time $t$, by combining previous studies to a common calibration following
@@
Figure 7 shows the $M_{*}$-SFR relation for the different redshift bins, color coded by their mean mass-weighted stellar ages. It can be seen that the main population of young, star forming galaxies (SFMS) matches with the empirical laws mostly for the low-$z$ subset, while a constant offset to higher SFRs with respect to the empirical formulation can be observed. This behaviour is similar to the overestimation bias measured with cigale on the simulations and likely has the same source, that is the simplified cigale SFH model that can heavily reduce the constraining power on related properties. Overall however, older, red and quiescent galaxies exhibit lower SFRs and lie below the SFMS, with intermediate galaxies where star formation has started quenching populating the green valley between the main clouds (Gallazzi et al., 2017; Kalinova et al., 2021).
@@
Now we separate the main galaxy populations by a cut of the specific star formation rate $\mathrm{sSFR}=\mathrm{SFR}/M_{*}$. The threshold is hereby set as
@@
as previously used by Weinmann et al. (2006), Salim et al. (2009), Lara-López et al. (2010), Annunziatella et al. (2014), Tortorelli et al. (2018). Applying this cut to the cigale output allows to plot the distributions of properties for these two main populations. Figure 8 shows again the distributions of estimated properties, now separated by their affiliation to either the red or blue galaxy population, respectively, as well as a histogram of the sSFR with the aforementioned cut.
@@ 5.2 SED fitting with Prospector
In this section, we analyze the PAUS narrow-bands and CFHTLS $ugriz$ observational data by running the Prospector MCMC estimation. We also plot the property distributions of the maximum a posteriori SED of the chains, which is displayed in Fig. 9. In this figure, the galaxies are already separated by the aforementioned sSFR cut into blue and red populations and shown with the respective cigale 2D distributions. We also plot the photometric redshift $z$, although this is not a free parameter of the fitting routine and just depicts the value from the PAUS catalog.
Figure 9: Distribution of MAP values of the selected properties as determined by Prospector: stellar mass, mass-weighted stellar age, SFR over 100 Myrs, stellar metallicity the optical depths of the ISM and BC dust components, all using both narrow-band and broad-band $ugriz$ data. We additionally show the redshift $z$. The two subplots depict the blue (left) and red (right) galaxy population after an sSFR cut, gray shows the respective populations from cigale.
@@
Overall, one can clearly see that the distributions appear largely different from the cigale results for some of the properties. For once, the mean age estimate for blue galaxies is much higher with Prospector, while it also finds best-fit values over a larger range of the age prior. Mass-weighted stellar ages there have values up to 10 Gyrs from Prospector-$\alpha$. This effect can be traced throughout the entire parameter space, where the Prospector overall property ranges strongly exceed the distribution widths from the previous cigale estimates, with especially the dust components and metallicity exhibiting a more widely spread posterior space. Still, the means of the 1D histograms are similiar between both codes, for instance for $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$.
@@
The galaxies however do not show a pronounced mass-bimodality, with both populations covering the full range of values as opposed to the cigale results. This could however be caused by the tight Prospector-$\alpha$ mass prior, which does not allow SED models with $M_{*}<10^{9}\,M_{\odot}$, resulting in a pile-up at the prior edge. Still, we observe a clear separation in mass-weighted ages and SFRs, however again not as clearly for the metallicity and $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$.
@@
The size of the posterior is here mainly related to the size of the prior, as cigale is not able to sample outside of its prior and thus cannot cover the same parameter ranges as Prospector-$\alpha$ due to the restrictive parameter space. We note however that increasing the grid values of cigale for example for the dust components did not yield better results. Hence, the differences are likely rooted in the sampling techniques and the different SPS models, mainly the star formation history complexity which, aside from age and SFR, especially influences metallicity and dust attenuation. The box-like shape of the Prospector metallicity distribution suggests that the accessible value range is either still not wide enough, that the prior is not informative enough (or the number of MCMC iterations is too low), making it hard to sample high-quality constraints, or that the photometry is not able to constrain this property accurately. Given the results for $Z_{*}$ on the simulation, we favor the latter explanation, given the inherent dust-age-metallicity degeneracy.
@@
The discrepancies between cigale and Prospector for $Z_{*}$, ages, $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$ are interesting to observe, as they indicate the strong dependence of SED fitting results on prior choices and SPS models, where just a more complex SFH model might strongly change the parameters of the best-fitting spectrum. The overall distributions of SFRs and stellar masses however are mostly similar. This is expected for stellar masses, however not naturally for the SFR, given that it is directly related to the SFH. Since the cigale parameter space was however chosen by analysis on Prospector-$\alpha$ simulations, a similar constraining power can be explained by the SPS model optimization process.
Given the similar measurement quality of the dust attenuation components for both codes on the simulations, the differences on the observations on these specific properties require further explanation. The distribution of Prospector-$\alpha$ priors, from which the simulated photometries where sampled, likely does not consistently cover the whole range of true parameters in the COSMOS field. cigale and Prospector will not access SPS models with parameters outside of the fixed grid or the prior, respectively, that we matched towards the simulation priors. As both codes handle the likelihood maximization very differently, discrepancies may naturally arise for some galaxies, that can be mostly attributed to the age-dust-metallicity degeneracy, which are the parameters where we observe the strongest differences. Besides, finding a low-$\chi^{2}$ fit is more complicated with many tightly spaced narrow-band fluxes, which may lead to inefficient exploration of the prior space or strong degeneracies.
@@
Broad-band SED fitting quality studies might have avoided such issues due to simpler template fitting approaches or smaller filter baselines, where the likelihood of a model SED is easier maximized given a low number of widely spaced filters. Previous narrow-band results from for example J-PAS also attributed differences in spectra and SED fits of narrow-band photometry to the signal-to-noise ratio and the age-metallicity degeneracy due to lack of corresponding indicators in the baseline (Mejía-Narváez et al., 2017). This impact might not be as important for photo-$z$ estimates, but requires further study on a broader range of SED codes to quantify the effect of SPS model choices. Still, we acknowledge that our Prospector results could demand a longer sampling process to cover the large high-dimensional prior space and thus might provide more similar results to cigale with longer chains or more walkers. On the other hand, the latter does not allow a flexible SFH and a more sophisticated sampling method, which could facilitate a better exploration of the parameter space. Such techniques, like MCMC, although infringe the code’s applicability to next generation surveys, where the amount of data requires faster techniques. In general however, one expects higher quality results from the MCMC fitting with a flexible non-parametric SFH prior, like in the Prospector-$\alpha$ model. Still, we do not know which of the results more closely matches the real values, as there exists no ground truth for observations and we thus cannot be certain if our priors tested on mocks are able to generate the real plethora of observed SEDs. A possible validation can be made through comparison with COSMOS2020, although we cannot be sure about the quality of the physical parameter estimates in the catalog either, due to the same limitations as with our setup.
@@
Figure 11: 1:1 relations comparing the results on the PAUS-COSMOS observations between both codes. On the $x$-axis we plot the cigale estimates on physical parameters, while the $y$-axis depicts the respective values from Prospector.
@@
Finally, we can also compare the results on a per galaxy basis, to show the real discrepancies between both codes below the population level analysis, which is depicted in Fig. 11. As discussed before, the stellar mass estimates are robust within 0.5 dex, aside from clipping at the prior edge with Prospector-$\alpha$. Similarly, SFRs are constrained well, which is not usually the case from photometry, suggesting a powerful hyper-parameter setup. While ages also seem constrained better than expected given the two different SFH approaches, a 0.5 dex level accuracy is large for estimates on a two orders of magnitude scale. Still, both codes find alike SFHs for a majority of the sample. The other properties on the other hand exhibit stronger discrepancies. Mainly the metallicity is not well constrained by the photometry as both codes produce extremely different values and distributions. For the dust parameters, cigale does not find best-fit models over the whole grid size, while Prospector-$\alpha$ covers a larger range of values, leading to matching values only close to the mean of the distribution. Given the varying $Z_{*}$ and $\langle t_{*}\rangle_{m}$, it is a natural consequence that also the third contribution to the aforementioned degeneracy does not produce a congruent relation between the two setups. To remove this issue and break the degeneracy, at least partly, in the future, sampling via a joint prior that is calibrated on a spectroscopically observed mass-metallicity relation (as done for the creation of simulated SEDs with the Gallazzi et al. 2005 prior), will be beneficial, although this is for now only possible for the Prospector code.
@@
Recent studies on the reliability of SED fitting techniques in general by Pacifici et al. (2023) tested the modelling uncertainties on different physical properties between different codes. They found that the stellar mass distribution is the most robust, with a higher uncertainty for SFH-related parameters (such as the SFR, especially at low-SFR) and dust properties, due to their correlation. This relates well to our findings, where mostly the dust optical depth and metallicity distributions can vary strongly between the two setups. Therefore, modelling choices, as well as caution during interpretation of SFH, dust and metallicity parameters without spectroscopic information, are of paramount importance for SED fitting.
@@
Aside from the statistical framework, the prior selection, the likelihood function and the dust-age-metallicity degeneracy, the two setups also differ in their stellar template libraries (MILES for Prospector and BC03 for cigale). Previous studies have tested the effect of varying spectral libraries with fixed fitting codes to test their respective contributions on physical property estimates. Coelho et al. (2009) and Dias et al. (2010) found that the choice of spectral templates (BC03, PEGASE-HR Le Borgne et al. 2004 and MILES) can have a significant impact on the age estimate in for example composite stellar populations of M32 or the Small Magellanic Cloud (SMC), without strongly affecting the overall metallicities derived from the integrated spectra. González Delgado & Cid Fernandes (2010) on the other hand found a stronger effect of the stellar library choice (MILES, GRANADA Martins et al. 2005, STELIB Le Borgne et al. 2003) on the metallicity, with typically 0.1 dex and 0.3 dex model-to-model dispersions for stellar ages and metallicities, respectively. This was however performed on stellar clusters, which can be characterized by a single age and do not require complex modelling a composite stellar populations over a wide range of ages. Recent work by Byrne & Stanway (2023) found a stronger impact of the stellar spectral library on stellar population synthesis, where especially line indices can vary strongly between different templates, leading also to deviations in age and metallicity. Extensive studies over the impact on the assumptions of SED models are also outlined in Siudek et al. (in prep), who found that using different SSPs leads to a median absolute bias on a 0.25 dex or 0.2 dex level for the stellar mass and the SFR, respectively. This difference can be related to the incorporation of incorporation of thermally pulsating asymptotic giant branch (TP-AGB) stars in Maraston et al. (2010) models, which are not included in the BC03 templates. Overall, this shows that the observed differences between the two setups are not only caused by the different code properties, but also by the basis of the applied SPS process itself. In App. B, we plot a comparison of simulated composite stellar populations from cigale and Prospector given different input SSPs and compare how such differences affect the simulated observed magnitudes in the PAUS and CFHTLS bands. We there show that even without influences from dust or gas models, the best-fit photometry from simulated spectra will vary at up to $\sim 0.2\,$mag level between our two setups for identical input physical parameters.

### v2

@@ 4.3 Results with Prospector
Figure 5: Hexbin plots showing the ground truth from the Prospector-$\alpha$ stellar population simulations on the x-axis versus the MAP value from t Prospector, for stellar mass, $M_{*}$, current SFR, mean mass-weighted stellar ages, $\langle t_{*}\rangle_{m}$, stellar metallicity, $Z_{*}$ and the optical depths of the two-component dust attenuation model, $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$.
@@
We find robust estimates for stellar masses and SFRs, which in this case is expected due to the same star formation histories used for the generation of the data and the forward-modeling of the fitted SED. Dust parameters are also recovered accurately on a 0.5 dex level for the majority of the sample galaxies, albeit with some objects where the birth cloud dust attenuation is not reliably recovered. The overall intrinsic scatter from the 1:1 relation apparent for every free parameter is a consequence of fits where the MCMC chain is not fully converged yet after the set number of iterations, as well as parameter degeneracies and the simulated uncertainties of the model photometry.
Our setup, however, struggles to estimate accurate stellar metallicity posteriors, mainly for older populations with high $Z_{*}$, where the estimated values seem to be mostly clipped to the edges of the uniform prior space. We attribute this to the prior differences during the simulation and the sampling phase, where we once chose a joint mass-metallicity prior and later a uniform distribution for sampling. Increasing the number of likelihood calls with more MCMC iterations or more walkers can help to alleviate this problem, as we restricted these hyper-parameters due to computational limitations. We show in Appendix A an example fit with an increased iteration count, which can lead to substantial shift of the best-fit parameters. Additionally, and most likely due to the age-metallicity degeneracy, $\langle t_{*}\rangle_{m}$ is not estimated well for a substantial subset of mock objects, where especially the ages of young stellar populations are overestimated by up to one order of magnitude. This then also influences the metallicity estimates and hinders an accurate measurement of both of these parameters without a joint prior.
@@
Additionally, we note that the reliance on only the MAP value of the posterior distribution is usually not always sufficient, mostly due to degeneracies of the parameters. Looking at some full posteriors from our estimates, one can identify multiple peaks with a high likelihood that where found exploring different regions of a joint dust-age-metallicity prior. We show an example of such a posterior distribution in Appendix A.
@@ 5 Stellar population properties
Using our selection of fitting hyper-parameters and cigale grid values, we show the results from both codes on the sample of 25 491 galaxies in the COSMOS field. We complete two runs for cigale on the catalog, once using CFHTLS $ugriz$ data only and once with additional 40 PAUS narrowband fluxes. For Prospector, we only run the code once with narrowband+broadband data, due to the required $\sim$ 500k CPU hours for one run on the full sample. We first show the overall distributions of our parameter estimation as well as the robustness of the inferred properties. We also perform a comparison between the two codes and the COSMOS2020 catalog to determine the predictive power of the PAUS narrowband filter set in relation to previous studies with broad filter baselines that also include narrowband photometries.
Figure 6: Distribution of MAP values of the selected properties as determined by cigale: stellar mass, mass-weighted stellar age, SFR over 100 Myrs, stellar metallicity and optical depths of ISM and BC dust components. Blue shows the results from using narrowband and broadband data, red depicts the properties from using CFHTLS $ugriz$ alone.
@@ 5.1 SED fitting with CIGALE
In Fig. 6, we show the distributions of the inferred physical properties obtained with cigale on our catalogue of 25 491 galaxies in the COSMOS field. The blue contours and histograms show the distributions resulting from running the SED fitting routine on the combined PAUS (narrowband) and CFHTLS (broadband) dataset, while the red contours were determined by using the broadband $ugriz$ data alone.
We first note that the stellar mass, which usually represents the most robust property inferred by SED fitting, is similarly constrained by either using only the broadband or the combination of narrowband and broadband data. The other properties, especially metallicity and dust component optical depths, show that with the addition of the narrowband, we can capture a broader range of physical values with respect to the sole use of the broadband. Since these physical properties are strongly correlated to distinct spectral features, the net effect of an increased photometric resolution manifests in the form of a more diverse distribution. Broadband filters alone tend to estimate metallicities that only slightly differ from $Z_{*}=Z_{\odot}$, as well as ages towards the high-end tail of the mass-weighted age distribution and low optical depths $\hat{\tau}_{\lambda,2;1}$ of the dust attenuation model. This can be interpreted as a result of the large impact of the prior choice and size on these parameters in the broadband-only case, resulting in a narrower posterior.
Nevertheless, the effect of the narrowband data is not as strong as expected, where most of the sample galaxies are still similarly fitted with reasonable property estimates from the underlying SPS forward-model using $ugriz$ data alone. However, we see that narrowbands do not lead to to a different region of the parameter space with respect to the use of broadband data only, which indicates that the model is performing in the same manner with either setup. Additionally, we note that the overall coverage of the prior ranges set by the cigale parameter space from Table 2 is largely diminished in the posterior distribution. This means that either the priors are too wide and the true observations do not cover the ranges set from the Prospector-$\alpha$ model, or the sparse grid leads to an incomplete sampling of the high-dimensional parameter space. This is facilitated by the degeneracies of certain model properties. We fill further explore this after analyzing the Prospector results and the comparison with COSMOS2020.
@@
with $S_{0}$, $M_{0}$, $\alpha$, and $\beta$ as empirical parameters that depend on redshift. Popesso et al. (2023) recently presented a new parametrization of the star forming main sequence as a function of time, $t$, by combining previous studies to a common calibration following:
@@
Figure 7 shows the $M_{*}$-SFR relation for the different redshift bins, color coded by their mean mass-weighted stellar ages. It can be seen that the main population of young, star forming galaxies (SFMS) matches with the empirical laws mostly for the low-$z$ subset, while a constant offset to higher SFRs with respect to the empirical formulation can be observed. This behavior is similar to the overestimation bias measured with cigale on the simulations and likely has the same source, that is the simplified cigale SFH model that can heavily reduce the constraining power on related properties. Overall however, older, red and quiescent galaxies exhibit lower SFRs and lie below the SFMS, with intermediate galaxies where star formation has started quenching populating the green valley between the main clouds (Gallazzi et al., 2017; Kalinova et al., 2021).
@@
Figure 9: Distribution of MAP values of the selected properties as determined by Prospector: stellar mass, mass-weighted stellar age, SFR over 100 Myrs, stellar metallicity the optical depths of the ISM and BC dust components, all using both narrowband and broadband $ugriz$ data. We additionally show the redshift $z$. The two subplots depict the blue and red galaxy population after an sSFR cut, gray shows the respective populations from cigale.
We go on to we separate the main galaxy populations by a cut of the specific star formation rate, $\mathrm{sSFR}=\mathrm{SFR}/M_{*}$. The threshold is hereby set as:
@@
as previously used by Weinmann et al. (2006), Salim et al. (2009), Lara-López et al. (2010), Annunziatella et al. (2014), Tortorelli et al. (2018). Applying this cut to the cigale output allows us to plot the distributions of properties for these two main populations. Figure 8 shows again the distributions of estimated properties, now separated by their affiliation to either the red or blue galaxy population, respectively, as well as a histogram of the sSFR with the aforementioned cut.
@@ 5.2 SED fitting with Prospector
In this section, we analyze the PAUS narrowbands and CFHTLS $ugriz$ observational data by running the Prospector MCMC estimation. We also plot the property distributions of the maximum a posteriori SED of the chains, which is displayed in Fig. 9. In this figure, the galaxies are already separated by the aforementioned sSFR cut into blue and red populations and shown with the respective cigale 2D distributions. We also plot the photometric redshift $z$, although this is not a free parameter of the fitting routine and just depicts the value from the PAUS catalog.
@@
Overall, we can clearly see that the distributions appear largely different from the cigale results for some of the properties. For once, the mean age estimate for blue galaxies is much higher with Prospector, while it also finds best-fit values over a larger range of the age prior. Mass-weighted stellar ages there have values up to 10 Gyrs from Prospector-$\alpha$. This effect can be traced throughout the entire parameter space, where the Prospector overall property ranges strongly exceed the distribution widths from the previous cigale estimates, with especially the dust components and metallicity exhibiting a more widely spread posterior space. Still, the means of the 1D histograms are similiar between both codes, for instance, for $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$.
@@
The galaxies however do not show a pronounced mass-bimodality, with both populations covering the full range of values as opposed to the cigale results. This could, however, be caused by the tight Prospector-$\alpha$ mass prior, which does not allow SED models with $M_{*}<10^{9}\,M_{\odot}$, resulting in a pile-up at the prior edge. Still, we observe a clear separation in mass-weighted ages and SFRs, however again not as clearly for the metallicity and $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$.
@@
The size of the posterior is here mainly related to the size of the prior, as cigale is not able to sample outside of its prior and thus cannot cover the same parameter ranges as Prospector-$\alpha$ due to the restrictive parameter space. We note however that increasing the grid values of cigale for example for the dust components did not yield better results. Hence, the differences are likely rooted in the sampling techniques and the different SPS models – mainly the star formation history complexity which, aside from age and SFR, especially influences metallicity and dust attenuation. The box-like shape of the Prospector metallicity distribution suggests that the accessible value range is either still not wide enough, that the prior is not informative enough (or the number of MCMC iterations is too low), making it hard to sample high-quality constraints, or that the photometry is not able to constrain this property accurately. Given the results for $Z_{*}$ on the simulation, we favor the latter explanation, given the inherent dust-age-metallicity degeneracy.
@@
The discrepancies between cigale and Prospector for $Z_{*}$, ages, $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$ are interesting to observe, as they indicate the strong dependence of SED fitting results on prior choices and SPS models, where just a more complex SFH model might strongly change the parameters of the best-fitting spectrum. The overall distributions of SFRs and stellar masses however are mostly similar. This is expected for stellar masses, however, this is not entirely natural for the SFR, given that it is directly related to the SFH. Since the cigale parameter space was however chosen by analysis on Prospector-$\alpha$ simulations, a similar constraining power can be explained by the SPS model optimization process.
Given the similar measurement quality of the dust attenuation components for both codes on the simulations, the differences on the observations on these specific properties require further explanation. The distribution of Prospector-$\alpha$ priors, from which the simulated photometries where sampled, likely does not consistently cover the whole range of true parameters in the COSMOS field. cigale and Prospector will not access SPS models with parameters outside of the fixed grid or the prior, respectively, that we matched towards the simulation priors. As both codes handle the likelihood maximization very differently, discrepancies may naturally arise for some galaxies, that can be mostly attributed to the age-dust-metallicity degeneracy, which are the parameters for which we observe the strongest differences. Besides, finding a low-$\chi^{2}$ fit is more complicated with many tightly spaced narrowband fluxes, which may lead to inefficient exploration of the prior space or strong degeneracies.
@@
Broadband SED fitting quality studies might have avoided such issues due to simpler template fitting approaches or smaller filter baselines, where the likelihood of a model SED is more easily maximized given a low number of widely spaced filters. Previous narrowband results from for example J-PAS also attributed differences in spectra and SED fits of narrowband photometry to the signal-to-noise ratio (S/N) and the age-metallicity degeneracy due to a lack of corresponding indicators in the baseline (Mejía-Narváez et al., 2017). The impact of this might not be as important for photo-$z$ estimates, but it would require further study on a broader range of SED codes to quantify the effect of SPS model choices. Still, we acknowledge that our Prospector results could demand a longer sampling process to cover the large high-dimensional prior space and thus might provide more similar results to cigale with longer chains or more walkers. On the other hand, the latter does not allow a flexible SFH and a more sophisticated sampling method, which could facilitate a better exploration of the parameter space. Such techniques, such as MCMC, although infringe the code’s applicability to next-generation surveys, where the amount of data requires faster techniques. In general however, one expects higher quality results from the MCMC fitting with a flexible non-parametric SFH prior, as in the Prospector-$\alpha$ model. Still, we do not know which of the results more closely matches the real values, as there exists no ground truth for observations and we thus cannot be certain if our priors tested on mocks are able to generate the real plethora of observed SEDs. A possible validation can be made through comparison with COSMOS2020, although we cannot be sure about the quality of the physical parameter estimates in the catalog either, due to the same limitations as with our setup.
@@
Figure 11: 1:1 relations to compare the results on the PAUS-COSMOS observations with both codes. On the $x$-axis we plot the cigale estimates on physical parameters, while the $y$-axis depicts the respective values from Prospector.
@@
Finally, we can also compare the results on a per galaxy basis to show the real discrepancies between both codes below the population level analysis (given in Fig. 11). As discussed earlier in this paper, the stellar mass estimates are robust within 0.5 dex, aside from clipping at the prior edge with Prospector-$\alpha$. Similarly, SFRs are constrained well, which is not usually the case from photometry, suggesting a powerful hyper-parameter setup. While ages also seem constrained better than expected given the two different SFH approaches, a 0.5 dex level accuracy is large for estimates on a two orders of magnitude scale. Still, both codes find alike SFHs for a majority of the sample. The other properties on the other hand exhibit stronger discrepancies. Mainly the metallicity is not well constrained by the photometry as both codes produce extremely different values and distributions. For the dust parameters, cigale does not find best-fit models over the whole grid size, while Prospector-$\alpha$ covers a larger range of values, leading to matching values only close to the mean of the distribution. Given the varying $Z_{*}$ and $\langle t_{*}\rangle_{m}$, it is a natural consequence that also the third contribution to the aforementioned degeneracy does not produce a congruent relation between the two setups. To remove this issue and break the degeneracy, at least partly, in the future, sampling via a joint prior that is calibrated on a spectroscopically observed mass-metallicity relation (as done for the creation of simulated SEDs with the Gallazzi et al. 2005 prior), will be beneficial, although this is for now only possible for the Prospector code.
@@
Recent studies on the reliability of SED fitting techniques in general by Pacifici et al. (2023) tested the modeling uncertainties on different physical properties between different codes. They found that the stellar mass distribution is the most robust, with a higher uncertainty for SFH-related parameters (such as the SFR, especially at low SFR) and dust properties, due to their correlation. This relates well to our findings, where mostly the dust optical depth and metallicity distributions can vary strongly between the two setups. Therefore, modeling choices, as well as caution during interpretation of SFH, dust and metallicity parameters without spectroscopic information, are of paramount importance for SED fitting.
@@
Aside from the statistical framework, the prior selection, the likelihood function and the dust-age-metallicity degeneracy, the two setups also differ in their stellar template libraries (MILES for Prospector and BC03 for cigale). Previous studies have tested the effect of varying spectral libraries with fixed fitting codes to test their respective contributions on physical property estimates. Coelho et al. (2009) and Dias et al. (2010) found that the choice of spectral templates (BC03, PEGASE-HR Le Borgne et al. 2004 and MILES) can have a significant impact on the age estimate in for example composite stellar populations of M32 or the Small Magellanic Cloud (SMC), without strongly affecting the overall metallicities derived from the integrated spectra. González Delgado & Cid Fernandes (2010) on the other hand found a stronger effect of the stellar library choice (MILES, GRANADA Martins et al. 2005, STELIB Le Borgne et al. 2003) on the metallicity, with typically 0.1 dex and 0.3 dex model-to-model dispersions for stellar ages and metallicities, respectively. This was however performed on stellar clusters, which can be characterized by a single age and do not require complex modeling a composite stellar populations over a wide range of ages. A recent work by Byrne & Stanway (2023) reported a stronger impact of the stellar spectral library on stellar population synthesis, where line indices, can particularly vary strongly between different templates, leading also to deviations in age and metallicity. Extensive studies over the impact on the assumptions of SED models are also outlined in Siudek et al. (in prep), who found that using different SSPs leads to a median absolute bias on a 0.25 dex or 0.2 dex level for the stellar mass and the SFR, respectively. This difference can be related to the incorporation of incorporation of thermally pulsating asymptotic giant branch (TP-AGB) stars in Maraston et al. (2010) models, which are not included in the BC03 templates. Overall, this shows that the observed differences between the two setups are not only caused by the different code properties, but also by the basis of the applied SPS process itself. In Appendix B, we plot a comparison of simulated composite stellar populations from cigale and Prospector given different input SSPs and compare how such differences affect the simulated observed magnitudes in the PAUS and CFHTLS bands. We there show that even without influences from dust or gas models, the best-fit photometry from simulated spectra will vary at up to $\sim 0.2\,$mag level between our two setups for identical input physical parameters.